OpenStudy (anonymous):

If there is an inverse relationship between x and y, and the constant of variation is 42, what is y when x=3? 1.9 2.39 3.14 4.126

OpenStudy (anonymous):

14

myininaya (myininaya):

y=k/x ----- y=42/x ----- y=42/3=14

OpenStudy (anonymous):

xy=42 3y=42 divide by 3 y=14